Stenson, Inc., imposes a payback cutoff of three years for its international investment projects. Assume the company has the following two projects available.Stenson, Inc., imposes a payback cutoff of three years for its international investment projects. Assume the company has the
Answer #1

Project Amount recovered in two year = 25000 + 32600 = 57600 Amount not recovered = 61000- 57600 - 3400 This remaining amount